# FlowSDF: Flow Matching for Medical Image Segmentation Using Distance Transforms

### v1

@@ Abstract
Medical image segmentation is a crucial task that relies on the ability to accurately identify and isolate regions of interest in medical images. Thereby, generative approaches allow to capture the statistical properties of segmentation masks that are dependent on the respective structures. In this work we propose FlowSDF, an image-guided conditional flow matching framework to represent the signed distance function (SDF) leading to an implicit distribution of segmentation masks. The advantage of leveraging the SDF is a more natural distortion when compared to that of binary masks. By learning a vector field that is directly related to the probability path of a conditional distribution of SDFs, we can accurately sample from the distribution of segmentation masks, allowing for the evaluation of statistical quantities. Thus, this probabilistic representation allows for the generation of uncertainty maps represented by the variance, which can aid in further analysis and enhance the predictive robustness. We qualitatively and quantitatively illustrate competitive performance of the proposed method on a public nuclei and gland segmentation data set, highlighting its utility in medical image segmentation applications.
@@ 1 Introduction
Medical image segmentation is the process of assigning each pixel in an image to a semantic class, which is crucial for extracting valuable insights for diagnosis and treatment planing. Traditionally, this task was handled by methods that have now been largely superseded by deep neural networks. These networks are particularly effective due to their ability to learn discriminatively from large data sets in an end-to-end manner (Ronneberger et al., 2015; Badrinarayanan et al., 2017; Wang et al., 2020). In addition to these discriminative models, generative models like generative adversarial networks (GANs) and denoising diffusion probabilistic models (DDPMs) have emerged for image segmentation with the advantage of learning the underlying statistics of segmentation masks conditioned on the respective images of interest (Iqbal et al., 2022; Xun et al., 2022; Amit et al., 2021). A very recently introduced generative approach is flow matching (Lipman et al., 2022; Liu et al., 2022), which uses an ordinary differential equation (ODE) to model the probability flow that transforms a data distribution into a simple, tractable prior distribution.
@@
To this end, we introduce FlowSDF, an approach that combines flow matching with the SDF for generative medical image segmentation. We build on prior work of (Bogensperger et al., 2023) where a variance-exploding (VE) stochastic differential equation (SDE) is employed to also model a conditional diffusion process to learn the statistics of segmentation masks given by the SDF. The generative diffusion model is improved upon by using an image-guided flow matching method (Lipman et al., 2022), which also learns a smooth implicit representation of segmentation masks conditioned on the corresponding images, see Figure 1. This integration allows the noise-injection perturbation process to canonically smooth the distribution of SDFs masks as further illustrated in Figure 4. Additionally, FlowSDF leverages its generative nature to produce multiple segmentations for a given input image, thereby allowing for segmentation uncertainty quantification. This also enhances the robustness and interpretability of our approach and provides valuable insights into the segmentation process. Summarized, our main contributions are:
@@
  • Further distance transforms are used to obtain a smoother, implicit representation of segmentation masks.
@@ 2.1 Medical Image Segmentation
In contrast, generative models for image segmentation such as GANs (Isola et al., 2017; Li et al., 2021) and score-based generative models (Baranchuk et al., 2021; Amit et al., 2021; Wolleb et al., 2022; Wu et al., 2022, 2024) have been utilized to focus on learning the data distributions of segmentation masks given input images. By capturing the underlying data distributions, these models provide more robust predictions, especially in scenarios with complex or unseen data. Although they offer statistical robustness, they generally require more complex training processes and tend to operate slower.
@@ 3.1 Image Segmentation using SDFs
Using SDFs the segmentation problem can be rephrased in the context of (signed) distances with respect to the object boundaries, where we consider each pixel $m_{ij}$ in the domain $\Omega$ with the object $\mathcal{S}$ to be segmented. The SDF map then contains for each $m_{ij}$ the distance to the closest boundary pixel $\partial\mathcal{S}$, where a negative/positive sign denotes the interior/outside of the object, respectively.
Mathematically, the SDF $\widetilde{m}\in\widetilde{\mathcal{M}}$ of a segmentation mask $m$ can thus be computed using the Euclidian distance function for each pixel $m_{ij}$, which additionally can be truncated at a threshold $\delta$ to consider only a span of pixels around the object boundary $\partial\mathcal{S}$. Moreover, the truncation helps to remain agnostic with respect to higher positive distances belonging to the background, as they should not impact the resulting segmentation. The implicit, truncated SDF $\phi(m_{ij})$ is then obtained as follows (also see (Osher et al., 2004)):
@@ 4.2 Training Details
During training, we utilized the Adam optimizer with a learning rate of 1e-4. The batch size was set to 16, and the models were trained for approximately 30k epochs for the first dataset and nearly 40k epochs for the second. Additionally, exponential moving average (EMA) with a decay factor of 0.9999 was applied throughout the training process. The detailed training procedure of image-guided flow matching is described in Algorithm 1.
@@ 4.3 Evaluation
After training we obtain a network $v_{\theta,\cdot}$ that regresses a vector field, capable of gradually transitioning $p(\widetilde{m}_{0})$ to $p(\widetilde{m}_{1}|x)$. Consequently, Algorithm 2 can be employed to synthesize new samples $\widetilde{m}_{1}$ given the conditioning images $x$ that have to be segmented. The ODEstep is implemented in practice using numerical integration provided by (Chen, 2018). The number of NFE is set to 4, as flow matching in general already yields remarkable results with only few NFEs, which is advantageous compared to standard DDPMs. Moreover, averaging over a number of $K$ predictions to obtain an estimate for the MMSE was considered, where the respective ablation that investigates the choice of $K$ can be found in Figure 9 for both data sets. For multiple noise injection runs $T$, we set $T=100$ to enable a fair comparison with the SDE sampler in (Amit et al., 2021).

### v2

@@ Abstract
Medical image segmentation plays an important role in accurately identifying and isolating regions of interest within medical images. Generative approaches are particularly effective in modeling the statistical properties of segmentation masks that are closely related to the respective structures. In this work we introduce FlowSDF, an image-guided conditional flow matching framework, designed to represent the signed distance function (SDF), and, in turn, to represent an implicit distribution of segmentation masks. The advantage of leveraging the SDF is a more natural distortion when compared to that of binary masks. Through the learning of a vector field associated with the probability path of conditional SDF distributions, our framework enables accurate sampling of segmentation masks and the computation of relevant statistical measures. This probabilistic approach also facilitates the generation of uncertainty maps represented by the variance, thereby supporting enhanced robustness in prediction and further analysis. We qualitatively and quantitatively illustrate competitive performance of the proposed method on a public nuclei and gland segmentation data set, highlighting its utility in medical image segmentation applications.
@@ 1 Introduction
Medical image segmentation is the process of assigning each pixel in an image to a semantic class, which is crucial for extracting valuable insights for diagnosis and treatment planing. footnotetext: Source code at https://github.com/leabogensperger/FlowSDF Traditionally, this task was handled by methods that have now been largely superseded by deep neural networks. These networks are particularly effective due to their ability to learn discriminatively from large data sets in an end-to-end manner (Ronneberger et al., 2015; Badrinarayanan et al., 2017; Wang et al., 2020). In addition to these discriminative models, generative models like generative adversarial networks (GANs) and denoising diffusion probabilistic models (DDPMs) have emerged for image segmentation with the advantage of learning the underlying statistics of segmentation masks conditioned on the respective images of interest (Iqbal et al., 2022; Xun et al., 2022; Amit et al., 2021). A very recently introduced generative approach is flow matching (Lipman et al., 2022; Liu et al., 2022), which uses an ordinary differential equation (ODE) to model the probability flow that transforms a data distribution into a simple, tractable prior distribution.
@@
To this end, we introduce FlowSDF, an approach that combines flow matching with the SDF for generative medical image segmentation. We build on prior work of (Bogensperger et al., 2023) where a variance-exploding (VE) stochastic differential equation (SDE) is employed to also model a conditional diffusion process to learn the statistics of segmentation masks given by the SDF. We improve upon our previous work in (Bogensperger et al., 2023) by using a generative flow matching framework (Lipman et al., 2022) in place of the generative diffusion model. Our proposed image-guided flow matching method also learns a smooth implicit representation of segmentation masks conditioned on the corresponding images, see Figure 1. This integration allows the noise-injection perturbation process to canonically smooth the distribution of SDFs masks as further illustrated in Figure 4. Additionally, FlowSDF leverages its generative nature to produce multiple segmentations for a given input image, thereby allowing for segmentation uncertainty quantification. This also enhances the robustness and interpretability of our approach and provides valuable insights into the segmentation process. Summarized, our main contributions are:
@@
  • Further, distance transforms are used to obtain a smoother, implicit representation of segmentation masks.
@@ 2.1 Medical Image Segmentation
Furthermore, when annotation data is available from multiple annotators to account for ambiguity in segmentation data, discriminative models can be transformed into probabilistic variants by incorporating stochastic information by sampling from a learned latent distribution using a conditional variational autoencoder (VAE) Kohl et al. (2018); Bhat et al. (2022). Several works build on this approach by utilizing the inherent hierarchical structure of U-Nets to integrate stochasticity Baumgartner et al. (2019); Kohl et al. (2019). Ambiguous data labels can also be exploited by generative models, such as diffusion models Rahman et al. (2023); however, this ambiguity is often not explicitly present in many data sets.
In contrast, generative models for image segmentation such as GANs (Isola et al., 2017; Li et al., 2021) and score-based generative models (Baranchuk et al., 2021; Amit et al., 2021; Wolleb et al., 2022; Wu et al., 2022, 2024) have been utilized to focus on learning the data distributions of segmentation masks given input images. While GANs are based on the principle of learning a direct mapping from noise to data through adversarial training, score-based diffusion models learn the log-gradient (score) of a data distribution by leveraging the idea of SDEs to gradually reverse a noise process. In our previous work, we use a score-based generative model where segmentation masks of corresponding medical images are represented implicitly by using SDFs (Bogensperger et al., 2023). By capturing the underlying data distributions, these models provide more robust predictions, especially in scenarios with complex or unseen data. Although they offer statistical robustness, they generally require more complex training processes and tend to operate slower.
@@ 3.1 Image Segmentation using SDFs
Using SDFs the segmentation problem can be rephrased in the context of Using SDFs, the segmentation problem can be reformulated in terms of signed distances relative to the object boundaries. For each pixel $m_{ij}$ in the domain $\Omega$ with the object $\mathcal{S}$ to be segmented, the SDF map assigns the distance to the nearest boundary pixel $\partial\mathcal{S}$. Additionally, the distance is signed, such that a negative sign indicates that the pixel is inside the object, while a positive sign indicates that it lies outside.
The SDF $\widetilde{m}\in\widetilde{\mathcal{M}}$ of a segmentation mask $m$ can be computed by applying the Euclidian distance function to each pixel $m_{ij}$. To focus on the region near the object boundary $\partial\mathcal{S}$, the distances can be truncated at a threshold $\delta$. This truncation also prevents distant background pixels from influencing the segmentation outcome, as large positive distances are ignored. The implicit, truncated SDF $\phi(m_{ij})$ is then given by (also see (Osher et al., 2004)):
@@ 4.2 Training Details
During training, we utilized the Adam optimizer with a learning rate of 1e-4. The batch size was set to 16, and the models were trained for approximately 30k epochs for the first data set and nearly 40k epochs for the second. The average training time per epoch was 5.4 seconds for the MoNuSeg data set and 15.5 seconds for the GlaS data set, using an NVIDIA Titan RTX GPU. Additionally, exponential moving average (EMA) with a decay factor of 0.9999 was applied throughout the training process. The detailed training procedure of image-guided flow matching is described in Algorithm 1.
@@ 4.3 Evaluation
After training we obtain a network $v_{\theta,\cdot}$ that regresses a vector field, capable of gradually transitioning $p(\widetilde{m}_{0})$ to $p(\widetilde{m}_{1}|x)$. Consequently, Algorithm 2 can be employed to synthesize new samples $\widetilde{m}_{1}$ given the conditioning images $x$ that have to be segmented. The ODEstep is implemented in practice using numerical integration provided by (Chen, 2018), requiring about 0.85 seconds to generate 4 samples with 4 NFEs. The number of NFE is set to 4, as flow matching in general already yields remarkable results with only few NFEs, which is advantageous compared to standard DDPMs. Moreover, averaging over a number of $K$ predictions to obtain an estimate for the MMSE was considered, where the respective ablation that investigates the choice of $K$ can be found in Figure 9 for both data sets. For multiple noise injection runs $T$, we set $T=100$ to enable a fair comparison with the SDE sampler in (Amit et al., 2021).